OpenStudy (anonymous):

What is the simplified form of the seventh root of x times the seventh root of x times the seventh root of x times the seventh root of x? 4 times the seventh root of x x to the four-sevenths power the seventh root of 4x x to the seven fourths power

OpenStudy (anonymous):

\[\sqrt[7]{x}\cdot\sqrt[7]{x}\cdot\sqrt[7]{x}\cdot\sqrt[7]{x}~~?\]

OpenStudy (anonymous):

\[\sqrt[n]{x}=x^{1/n}~~\Rightarrow~~\sqrt[7]{x}=x^{1/7}\] \[x^{a}x^{b}=x^{a+b}~~\Rightarrow~~x^{1/7}\cdot x^{1/7}\cdot x^{1/7}\cdot x^{1/7}=x^{4/7}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw